OpenStudy (anonymous):

Solve: −2(x − 5) = −4 7 −7 negative 1 over 2 −3

OpenStudy (anonymous):

will give medal

OpenStudy (anonymous):

7

OpenStudy (anonymous):

because if u have -2(x-5) you have to have x be positive othersie you anser will be negative.

OpenStudy (anonymous):

**Distributive property. a(b + c) = ab + ac -2(x - 5) = -4 -2x + 10 = -4 Solve for x. -2x + 10 = -4 Subtract 10. -2x = -14 Divide by -2. x = 7

OpenStudy (anonymous):

2 neg = a pos

OpenStudy (anonymous):

Thanks both of yall.... i suck at algebra lol :)

OpenStudy (jiteshmeghwal9):

distribute -2 on left side -2(x)-(-2)(5)=-4 -2x+10=-4 10+4=2x x=10+4/2=?

OpenStudy (anonymous):

yw (: I did too, but after a while it kinda sticks lol

OpenStudy (anonymous):

yes:)

OpenStudy (anonymous):

yea well it wouldnt be so bad if i wasnt doing virtual school... i hate it so much

OpenStudy (anonymous):

i have the same view lol

OpenStudy (anonymous):

well im playing high school baseball but my mom doesnt want me in public school so i have to homeschooled to play on the team lol but i can go to a school that doesnt have a baseball team ... prob going nxt year hopefully

OpenStudy (anonymous):

yeah im going to public nxt years too